Two football players are pushing a 60 kg blocking sled across the field at a constant speed of 2.5 m/s. The coefficient of kinetic friction between the grass and the sled is 0.30. Once they stop pushing, how far will the sled slide before coming to rest?
Answer #1

this is a question of conservation of energy

KE imparted by the players will be lost into overcoming the friction

0.5mv^2 = umg*S

2.5^2 = 2*9.8*0.3*S

s = 1.06 m